OpenStudy (anonymous):

write an equation to represent the given relationship among integers. the sum of three consecutive integers is 69.

OpenStudy (ganpat):

let 1st integer be x .. so the consecutive integers would be, x, x+1 and x+2.. can u proceed now ?

OpenStudy (anonymous):

not really confused

OpenStudy (ganpat):

ok.. they said 3 consecutive numbers... they can be any... and consecutive means, 1, 2 & 3 or 4, 5 & 6... so we r taking three numbers as x, x +1 and x +2 .. as we do not know, what the numbers are and we intend to find that... so add them and equate to 69..

OpenStudy (ganpat):

gotta ??

OpenStudy (ganpat):

x, x +1 and x +2 = 69.. can u solve further ??

OpenStudy (ganpat):

x + x + 1 + x +2 = 69 3x + 3 = 69.. divide throughout by 3.. x + 1 = 13.. So x = 12... and the three numbers r , 12, 13 & 14... understood ??